OpenStudy (anonymous):

David an Fiona have a house painting business. It takes Fiona 3 days to paint a certain house. David could paint the same house in 4 days. How long would it take them to paint the house if David and Fiona worked together?

OpenStudy (anonymous):

@agent0smith Do you know how to o this?

OpenStudy (agent0smith):

this is the way i do them\[\Large \frac{ 1 }{ t } = \frac{ 1 }{ 4 }+\frac{ 1 }{3 }\]solve for t

OpenStudy (anonymous):

that's what I did but I got the answer 7/2 as the final answer.but that's wrong. Can you go throught the steps please?

OpenStudy (anonymous):

sorry 7/12

OpenStudy (agent0smith):

\[\Large \frac{ 1 }{ t } = \frac{ 7 }{ 12 }\] that doesn't mean the answer is 7/12

OpenStudy (anonymous):

oh would I cross Multiply now?

OpenStudy (agent0smith):

Yes. Orrrr just flip both sides since they are both fractions. 1/t becomes t/1.

OpenStudy (anonymous):

Alright, thankyou! :) Can you help me with one more?